OpenStudy (anonymous):

What is -2 1/2 divided by 6 ?

OpenStudy (anonymous):

@pooja195

OpenStudy (alexandervonhumboldt2):

-2 1/2*6=-5/2*1/6=-5/12

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

go to this website

OpenStudy (anonymous):

did it help